OpenStudy (anonymous):

find dy/dx from first principles given y=4-x

OpenStudy (cwrw238):

dy /dx is limit as h --> 0 of f(x + h) - f(x) ------------ h

OpenStudy (anonymous):

so where does y=4-x go to?

OpenStudy (kainui):

y is a function of x, so you can rewrite y=4-x as f(x)=4-x

OpenStudy (anonymous):

thankyou